User name jpchen
Log entry time 19:10:09 on February15,2002
Entry number 82834
The cryotarget behave strangely: there are large temperature difference
between the sensors before and after the target cell (about 4 degrees!).
With PID, high power heater was jumping from 0 to maximum (400 W)
while tempearture also fluctuate about +- 0.5 degree around 19 K for CT119.
Seems the loop circulation is very poor. Since the techometer gives reading
of 60 Hz, probably one fan is working fine, and another one might be not working?
I changed the loop set point from 19 to 18K (at 18 K, the average temperature
is still about (18+22)/2=20K.) and changed the PID values from (600, 40, 0)
to (60, 10, 0). It is now a little more stable (fluctuate at +-0.1 K) and the heater
responds much slower. The temperature readings with 40 uA on the optics target are:
CT119=18.00K, CT118=18.05K, CT115=22.12K, CT116=22.0K.
Will be useful to record the values when beam is off.